# A Survey on Diffusion Models for Time Series and Spatio-Temporal Data

### v3

@@ 2.3.4 Improvements with Diffusion Model and Its Variants
While diffusion models have generally yielded satisfactory results in generation and various tasks, practical applications reveal certain limitations. These include the slow iterative sampling process and the computational complexity arising from high-dimensional input, which affects efficiency. Furthermore, there are concerns regarding their generalization performance across different distributions and the challenges in integrating them with other generative models. Next, we will explore the variants of diffusion models from efficiency and performance improvement, as shown in Fig. 5, highlighting the modifications and optimizations they bring in comparison to original diffusion models.
@@
Efficiency Enhancement. (1) Forward Processes Improvement shifted towards leveraging other physical phenomena to enhance model efficiency and robustness. For example, inspired by electric field dynamics, PFGM [211] and its extension PFGM++ [212] [213] were proposed. They are guided distributions along electric field lines and employing augmented dimensions for improved performance. Moreover, innovations like Cold Diffusion explore the use of image transformations as a forward process [214]. Also, some improved Gaussian perturbation kernel methods were presented [215]. (2) Reverse Processes Improvement usually reduces the number of generation steps or uses lightweight models to enhance efficiency. Training-free sampling methods offer a paradigm where the acceleration of the sampling process does not necessitate retraining and optimize the trajectory from noise to data distributions, such as ODE-based methods DDIM [12], its extensions gDDIM [216], PNDM [145], EDM [215], DEIS [217], DPM-Solver [146], etc, and SDE-based methods through advanced techniques like restart sampling [218, 219]. On the other hand, some methods apply knowledge distillation to create an efficient and smaller network by transferring insights from large, intricate teacher models to simpler student ones [220, 221]. (3) Integration with Other Models is a common trick to enhance efficiency, such as combining with VAE [224, 225] and GAN [222, 223]. Additionally, there are methods that use the latent space as an input for the diffusion model, which can improve efficiency by reducing the input dimensionality [225, 7, 227, 228]. (4) Scheduler Functions improve efficiency by optimizing the reverse process, enabling faster convergence and reducing the number of required iterations. Common improved schedulers and their algorithms include CMS [221], DDIM [12], IDDPM [3], DEIS [217], DPM-Solvers [146, 229], Euler and Heun scheduler [215], LCM [230], RePaintScheduler [231], TCD [232], UniPC [233], and VQD [234].
@@
Performance Enhancement. (1) Model Architecture. Traditional DDPM uses the UNet architecture for its efficiency in reducing computational costs by leveraging a detailed feature space. UNet-based Innovations included normalization [236, 220], and different attentions with position encoding  [217, 237, 235]. Recently, to improve performance, transformer-based architecture is being used more often [152]. Reverse noise prediction incorporates adaptations to accommodate temporal and conditional inputs. Common methods include ViT [239], SwinTransformer [241, 240], DiT [242], and SiT [243]. (2) Training Schedules involve advanced decoding strategies, mainly classified into optimizing the diffusion stages and innovative projection methods. Optimization strategies focus on adjusting the diffusion steps to fine-tune the model performance [222, 224]. Meanwhile, projection techniques investigate various diffusion processes, like using linear distortions and different kernels, to increase model versatility [244, 245]. Moreover, innovations in diffusion models tweak noise adjustment for better output and quicker alignment [235, 246, 3]. Also, there are some optimization methods focusing on loss and matching [247, 248]. (3) Data Bridging is designed to address the limitations of generating arbitrary Gaussian distributions with complex distributions. Research has innovated with SDE/ODE principles to bridge gaps between distributions. Techniques like $\alpha$-blending create deterministic paths, using diffusion models for Gaussian-related cases [249]. Rectified Flow and other methods introduce optimizations and explore ODE creation between distributions [226, 250]. Additionally, leveraging the Schrödinger Bridge concept or Gaussian distributions as intermediaries offers new avenues in distribution transportation [251, 252]. (4) Conditional and Guidance Strategies are crucial for diffusion models, directing generation and enhancing output relevance and quality in response to specific conditions. Related works have been discussed in Sec. 2.3.3.
@@
Currently, research institutions, such as OpenAI (DALL-E series, GLIDE, SORA) and Stability AI (Stable Diffusion), have launched numerous outstanding diffusion models, which have exploded in popularity in both academic research and applications. For instance, Stable Diffusion is a state-of-the-art diffusion model known for its ability to generate high-quality, detailed images from textual descriptions, offering a powerful tool for creative and generative tasks [7]. ControlNet is a framework designed to enhance control over the attributes and structure of generated images, enabling precise manipulation of visual elements in generative models [199]. The consistency model achieves state-of-the-art performance without the slow iterative process of traditional diffusion models and offers capabilities like zero-shot editing and efficient training options [221].
2.4 Diffusion Model vs. Other Generative Models
Figure 6: Different types of generative models. (a) VAE is a probabilistic generative model based on encoding data into a latent space and decoding from this space for generation. (b) GAN generates new data similar to the training data by pitting two neural networks, a generator, and a discriminator, against each other in a game-like scenario. (c) Flow-based model generates data using invertible transformations, enabling precise computation of the probability density function of the data. (d) The diffusion model learns the underlying distribution of data by gradually introducing and then removing noise, enabling the generation of high-quality and diverse data samples.
@@
In addition to the diffusion model, there are many classical generative models, we take the most widely used variational autoencoders (VAEs), generative adversarial networks (GANs), and flow-based generative models as examples to introduce them and analyse their advantages, disadvantages and differences with the diffusion model. The workflow of these models is shown in Fig. 6.
VAE is a probabilistic generative model that encodes input data into a latent space $z$ and decodes from that latent space to generate data [186], as shown in Fig. 6(a). VAEs aim to maximize the lower bound of the log-likelihood of the data, known as the Evidence Lower Bound (ELBO).
$\mathsf{ELBO}=\mathbb{E}_{q_{\phi}(z|x)}[\mathsf{log}p_{\theta}(x|z)]+\mathsf{KL}[q_{\phi}(z|x)||p_{\theta}(z)],$ (17)
where $q_{\phi}(z|x)$ is the posterior distribution of the latent space output by the encoder and $p_{\theta}(x|z)$ is the conditional distribution of the data given the latent variable $z$.
GAN consists of a generator $G(z)$ and a discriminator $D(x,x^{\prime})$, as shown in Fig. 6(b). The generator tries to produce samples as close to the real data $x$ as possible, while the discriminator tries to distinguish between real data $x$ and generated data $x^{\prime}$. The training of GANs involves a zero-sum game, continually optimizing both the generator and the discriminator [187]. The objective function can be summarised as follows:
$\mathsf{min}_{G}\mathsf{max}_{D}V(D,G)=\mathbb{E}_{x\sim p_{data(x)}}[\mathsf{log}D(x,x^{\prime})]$ (18)
$+\mathbb{E}_{z\sim p_{z}(z)}[\mathsf{log}(1-D((G(z)))].$
Flow-based generative model (normalizing flows) transform data into a simpler distribution (e.g., Gaussian) through a sequence of reversible transformations, as shown in Fig. 6(c), ensuring precise likelihood evaluation due to the transformations being invertible [188]. The flow-based transformation formula is:
$z=f(x);\quad x^{\prime}=f^{-1}(z),$ (19)
where $x$ and $z$ are the points in the data space and latent space, respectively. $f$ is an invertible function. The log-likelihood of the transformation is given by:
$\mathsf{log}p(x)=\mathsf{log}p_{z}(f(x))+\mathsf{log}\left|\mathsf{det}(\frac{df}{dx})\right|.$ (20)
@@
Actually, different generative models each have their own unique strengths and limitations, making them suitable for specific applications. VAE is appreciated for its simplicity, stability, and clear theoretical underpinnings, yet it tends to produce lower-quality data and has limited expressiveness in its latent spaces. GAN is renowned for having powerful generative capabilities but is notoriously difficult to train and prone to instability and mode collapse. The flow-based model provides precise likelihood estimation and facilitates high-quality generation, but it demands significant computational resources and involves complex model designs. Diffusion models excel in generating high-quality and detailed output, offering flexibility and solid probabilistic foundations, but suffer from long training times and high computational costs. However, many variants of diffusion models and improvements have mitigated these issues, as described in Sec. 2.3.4. Currently, the diffusion model is the most popular generative model.
3 Overview and Categorization
@@
TABLE I: Summary and main papers of the diffusion models for time series and spatio-temporal data modeling. Red indicates univariate time series, blue-violet is multivariate time series, and yellow shows spatio-temporal data.
@@
WaveGrad [20] Univariate DDPM Generation Audio Google ICLR 2021
@@
DiffWave [91] Univariate DDPM Generation Audio UCSD ICLR 2021
D-Va [47] Univariate DDPM Forecasting Finance NUS CIKM 2023
DiffLoad [28] Univariate DDPM Forecasting Electricity HKU ArXiv 2023
TDSTF [289] Univariate DDPM Forecasting Healthcare U of A ArXiv 2023
DiffuASR [23] Univariate DDPM Generation Recommendation XJTU CIKM 2023
DiffRec [132] Univariate DDPM Generation Recommendation NUS SIGIR 2023
RecFusion [130] Univariate DDPM Generation Recommendation UvA ArXiv 2023
DiffRec [134] Univariate DDPM Generation Recommendation SCU ArXiv 2023
DiffuRec [95] Univariate DDPM Generation Recommendation WHU ArXiv 2023
PDRec [135] Univariate DDPM Generation Recommendation SDU ArXiv 2024
@@
TimeGrad [13] Multivariate DDPM Forecasting General Zalando ICML 2021
CARD [98] Multivariate DDPM Classification General UT-Austin NeurIPS 2022
BVAE [86] Multivariate DDPM Forecasting General Baidu NeurIPS 2023
TSDiff [164] Multivariate DDPM Forecasting General Amazon NeurIPS 2023
Kuo et al. [19] Multivariate DDPM Generation Healthcare UNSW ArXiv 2023
TS-Diffusion [278] Multivariate DDPM Generation General Cambridge ArXiv 2023
@@
SSSD [275] Multivariate DDPM Imputation General Oldenburg TMLR 2023
DA-TASWDM [276] Multivariate DDPM Imputation Healthcare HKBU CIKM 2023
Pintilie et al. [259] Multivariate DDPM Anomaly Detection General Bucharest ICDM 2023
DDMT [99] Multivariate DDPM Anomaly Detection General FNU ArXiv 2023
D3A-TS [101] Multivariate DDPM Imputation&Generation General Seville ArXiv 2023
TSDM [103] Multivariate DDPM Generation Environment HIT ArXiv 2023
TimeDDPM [102] Multivariate DDPM Generation Environment ZUST IEEE Sens. J. 2023
DiffEEG [18] Multivariate DDPM Forecasting Healthcare USTC ArXiv 2023
Diff-E [89] Multivariate DDPM Classification Healthcare Korea University Interspeech 2023
Tosato et al. [119] Multivariate DDPM Classification Healthcare Tilburg Synapsium 2023
DiffCharge [96] Multivariate DDPM Generation Electricity HKUST ArXiv 2023
@@
Yang et al. [238] Multivariate DDPM Imputation&Anomaly Detection AIOps Microsoft ESEC/FSE 2023
@@
AnoDDPM [46] Multivariate DDPM Anomaly Detection General Beihang IEEE Sens. J. 2024
DiffShape [277] Multivariate DDPM Classification General SCUT AAAI 2024
@@
STPP [70] Spatio-temporal DDPM Forecasting General Event THU KDD 2023
@@
ERDiff [25] Spatio-temporal DDPM Alignment General GIT NeurIPS 2023
DVGNN [59] Spatio-temporal DDPM Forecasting Transportation Halmstad ArXiv 2023
Yun et al. [69] Spatio-temporal DDPM Imputation General KAIST OpenReview 2023
DiffTAD [83] Spatio-temporal DDPM Anomaly Detection Transportation Xidian KNOWL-BASED SYST 2024
@@
SpecSTG [88] Spatio-temporal DDPM Forecasting Transportation USYD ArXiv 2024
@@
DST-DDPM [78] Spatio-temporal DDPM Forecasting Environment HKUST ENVIRON RES 2024
Westny et al. [80] Spatio-temporal DDPM Forecasting Transportation LiU ArXiv 2024
@@
UTD-PTP [79] Spatio-temporal DDPM Forecasting Transportation BIT IEEE Sens. J. 2024
@@
SGMSE [284] Univariate Score-based Generation Audio UHH Interspeech 2022
@@
DeScoD-ECG [57] Univariate Score-based Denoising Healthcare U of A IEEE J BIOMED HEALTH 2023
ScoreGrad [71] Multivariate Score-based Forecasting General BIT ArXiv 2021
Bilo et al. [104] Multivariate Score-based Forecasting General TUM ICML 2023
@@
StoRM [105] Multivariate Score-based Denoising&Generation Audio UHH IEEE-ACM T AUDIO SPE 2023
Lay et al. [178] Multivariate Score-based Denoising&Generation Audio UHH Interspeech 2023
@@
Risk-sensitive SDE [280] Multivariate Score-based Generation General Cambridge ArXiv 2024
@@
TimeADDM [282] Multivariate Score-based Anomaly Detection General HFUT ICASSP 2024
@@
Sasdim [72] Spatio-temporal Score-based Imputation General CSU ArXiv 2023
@@
Dyffusion [285] Spatio-temporal Score-based Forecasting General UCSD NeurIPS 2023
@@
Lu et al. [175] Univariate Conditional Generation Audio CMU ICASSP 2022
Wang et al. [142] Univariate Conditional Generation Electricity U of Macau ArXiv 2023
PulseDiff [144] Univariate Conditional Imputation Healthcare IC ArXiv 2023
DCDR [131] Univariate Conditional Generation Recommendation Kuaishou ArXiv 2023
Wang et al. [36] Univariate Conditional Generation Recommendation UIC ArXiv 2023
GiffCF [38] Univariate Conditional Generation Recommendation USTC ArXiv 2023
DR-DiffuSE [176] Univariate Conditional Generation Audio UESTC AAAI 2023
Dose [177] Univariate Conditional Generation Audio UESTC NeurIPS 2023
CRA-DIFFUSE [179] Univariate Conditional Generation Audio XJU ICME 2023
@@
DiffsFormer [52] Univariate Conditional Imputation&Generation Finance USTC ArXiv 2024
@@
DreamRec [167] Univariate Conditional Generation Recommendation USTC NeurIPS 2024
@@
CSDI [56] Multivariate Conditional Imputation General Stanford NeurIPS 2021
SF-DM [169] Multivariate Conditional Classification Manufacture Aalto UBICOMP 2023
@@
TimeDiff [34] Multivariate Conditional Forecasting General HKUST ICML 2023
@@
DiffAD [260] Multivariate Conditional Anomaly Detection General HENU KDD 2023
@@
D3[261] Multivariate Conditional Anomaly Detection General BUPT NeurIPS 2023
CLDM [107] Multivariate Conditional Generation Energy NCEPU IEEE T SUSTAIN ENERG 2023
MIDM [262] Multivariate Conditional Imputation General USTC KDD 2023
MEDiC [110] Multivariate Conditional Imputation Healthcare IIT NeurIPS 2023
VGCDM [263] Multivariate Conditional Anomaly Detection Electricity XJTU ArXiv 2023
DiffPLF [143] Multivariate Conditional Forecasting Electricity HKUST(GZ) ArXiv 2024
Time Weaver [182] Multivariate Conditional Generation General UT-Austin ArXiv 2024
Fu et al. [266] Multivariate Conditional Generation Electricity NUS ArXiv 2024
@@
ImDiffusion [257] Multivariate Conditional Anomaly Detection General PKU VLDB 2024
Wang [264] Multivariate Conditional Forecasting Business USTC ICASSP 2024
@@
DiffDA [26] Multivariate Conditional Forecasting Climate ETH Zurich ArXiv 2024
BioDiffusion [51] Multivariate Conditional Generation Healthcare TSU ArXiv 2024
@@
DiffSTOCK [41] Multivariate Conditional Forecasting Finance Purdue ICASSP 2024
RF-Diffusion [44] Multivariate Conditional Generation Network THU MobiCom 2024
Klein et al. [42] Multivariate Conditional Generation Healthcare Radboud ArXiv 2024
@@
DiffSTG [43] Spatio-temporal Conditional Forecasting General BJTU SIGSPATIAL 2023
@@
PriSTI [111] Spatio-temporal Conditional Imputation General Beihang ICDE 2023
@@
DiffUFlow [35] Spatio-temporal Conditional Forecasting Transportation CSU CIKM 2023
@@
DiffTraj [81] Spatio-temporal Conditional Generation Transportation SUSTech NeurIPS 2023
@@
ControlTraj [290] Spatio-temporal Conditional Generation Transportation HKUST(GZ) arXiv 2024
USTD [170] Spatio-temporal Conditional Forecasting General NUS ArXiv 2023
IDM [184] Spatio-temporal Conditional Forecasting Transportation ZJU ArXiv 2024
@@
Diff-RNTraj [17] Spatio-temporal Conditional Generation Transportation BJTU ArXiv 2024
Evans et al. [185] Univariate LDM Generation Audio Stability AI ArXiv 2024
LDT [253] Multivariate LDM Forecasting General NTU AAAI 2024
Aristimunha et al. [128] Multivariate LDM Generation Healthcare UPSaclay NeurIPS 2023
@@
TSDM [269] Multivariate DDIM Imputation&Anomaly Detection Electricity HUST ArXiv 2023
@@
LADM [265] Spatio-temporal LDM Forecasting Transportation XJU IEEE T INSTRUM MEAS 2024
{forest}
edge=thick forked edges, for tree= fill=level0!80, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8em,font=,fill=pink!20,, where level=2text width=9.5em,font=,fill=red!10,, where level=3text width=8em,font=,fill=blue!10,, where level=4text width=10em,font=,fill=gray!20,, where level=5text width=5em,font=,, [ Diffusion Models for

### v5

@@ 2.3.4. Improvements with Diffusion Model and Its Variants
While diffusion models have yielded satisfactory results in various tasks, practical applications reveal limitations such as slow iterative sampling and computational complexity due to high-dimensional input. There are also concerns regarding generalization performance and integration with other generative models. Next, we will explore the variants of diffusion models from efficiency and performance improvement, highlighting the modifications and optimizations they bring in comparison to the original diffusion models.
@@
Efficiency Enhancement. (1) Forward Processes Improvement: Several innovations aim to enhance model efficiency and robustness by using physical phenomena. PFGM (Xu et al., 2022a) and its extension PFGM++ (Xu et al., 2023b) and (Liu et al., 2023d) use electric field dynamics and augmented dimensions for improved performance. Cold Diffusion uses image transformations as a forward process (Bansal et al., 2022), and improved Gaussian perturbation kernel methods were presented (Karras et al., 2022). (2) Reverse Processes Improvement: Methods to reduce generation steps or use lightweight models include training-free sampling methods that optimize the trajectory from noise to data distributions, such as DDIM (Song et al., 2020a), gDDIM (Zhang et al., 2022), PNDM (Liu et al., 2022c), EDM (Karras et al., 2022), DEIS (Zhang and Chen, 2022), and DPM-Solver (Lu et al., 2022b). SDE-based methods use techniques like restart sampling (Jolicoeur-Martineau et al., 2021; Xu et al., 2023a). Knowledge distillation is also used by transferring insights from larger models to simpler ones (Berthelot et al., 2023; Song et al., 2023). (3) Integration with Other Models: Combining diffusion models with VAE (Lyu et al., 2022; Pandey et al., 2022) and GAN (Zheng et al., 2022; Watson et al., 2021) improves efficiency. Using latent space as input can reduce dimensionality and enhance efficiency (Pandey et al., 2022; Rombach et al., 2022; Vahdat et al., 2021; Zhang et al., 2023b). (4) Scheduler Functions: Optimizing the reverse process with improved schedulers such as CMS (Song et al., 2023), DDIM (Song et al., 2020a), IDDPM (Nichol and Dhariwal, 2021), DEIS (Zhang and Chen, 2022), DPM-Solvers (Lu et al., 2022b, c), Euler and Heun scheduler (Karras et al., 2022), LCM (Luo et al., 2023), RePaintScheduler (Lugmayr et al., 2022), TCD (Zheng et al., 2024), UniPC (Zhao et al., 2024a), and VQD (Gu et al., 2022) enables faster convergence and reduces iterations.
@@
Performance Enhancement. (1) Model Architecture: Traditional DDPM uses the U-Net architecture for efficiency. Innovations include normalization (Kawar et al., 2022; Berthelot et al., 2023) and attention mechanisms with position encoding (Zhang and Chen, 2022; Rissanen et al., 2022; Kingma et al., 2021). Transformer-based architectures are increasingly used to improve performance (Vaswani et al., 2017), including ViT (Dosovitskiy et al., 2020), SwinTransformer (Liu et al., 2021a, 2022b), DiT (Peebles and Xie, 2023), and SiT (Ma et al., 2024a). (2) Training Schedules: Advanced decoding strategies involve optimizing diffusion stages and innovative projection methods. Optimization strategies adjust the diffusion steps (Zheng et al., 2022; Lyu et al., 2022), while projection techniques use various diffusion processes to increase versatility (Hoogeboom and Salimans, 2022; Daras et al., 2022). Innovations in noise adjustment (Kingma et al., 2021; Kong and Ping, 2021; Nichol and Dhariwal, 2021) and optimization methods focusing on loss and matching (Song et al., 2021; Huang et al., 2021) enhance model performance. (3) Data Bridging: To address generating arbitrary Gaussian distributions with complex distributions, research uses SDE/ODE principles. Techniques like $\alpha$-blending (Heitz et al., 2023) and Rectified Flow (Liu et al., 2022a; Albergo and Vanden-Eijnden, 2022) explore ODE creation between distributions. The Schrödinger Bridge concept offers new avenues in distribution transportation (Su et al., 2022; Liu et al., 2023e). (4) Conditional and Guidance Strategies: These strategies are crucial for directing generation and enhancing output relevance and quality in response to specific conditions, as discussed in Sec 2.3.3.
@@
Currently, institutions like OpenAI and Stability AI have launched numerous outstanding diffusion models. For instance, Stable Diffusion generates high-quality images from textual descriptions (Rombach et al., 2022), and ControlNet enhances control over generated images (Zhang et al., 2023c). The consistency model offers state-of-the-art performance without the slow iterative process of traditional models (Song et al., 2023).
2.4. Diffusion Model vs Other Generative Models
Figure 5. Different types of generative models.
@@ 2.4. Diffusion Model vs Other Generative Models
In addition to the diffusion model, there are many classical generative models. We take the most widely used variational autoencoders (VAEs), generative adversarial networks (GANs), and flow-based generative models as examples to introduce them and analyse their advantages, disadvantages and differences with the diffusion model. The workflow of these models is shown in Figure 5 .
VAE is a probabilistic generative model that encodes input data into a latent space $z$ and decodes from that latent space to generate data (Kingma and Welling, 2013), as shown in Figure 5 (a). VAEs aim to maximize the lower bound of the log-likelihood of the data, known as the Evidence Lower Bound (ELBO). GAN consists of a generator $G(z)$ and a discriminator $D(x,x^{\prime})$, as shown in Figure 5 (b). The generator tries to produce samples as close to real data $x$ as possible, while the discriminator tries to distinguish between real data $x$ and generated data $x^{\prime}$. The training of GANs involves a zero-sum game, continually optimizing both the generator and the discriminator (Goodfellow et al., 2014). Flow-based generative model (normalizing flows) transform data into a simpler distribution (e.g., Gaussian) through a sequence of reversible transformations, as shown in Figure 5 (c), ensuring precise likelihood evaluation due to the transformations being invertible (Rezende and Mohamed, 2015).
@@
Different generative models have their unique strengths and limitations, making them suitable for specific applications. VAE is appreciated for its simplicity, stability, and clear theoretical foundation, but tends to produce lower-quality data and has limited expressiveness in its latent spaces. GAN is known for its powerful generative capabilities but is notoriously difficult to train and is prone to instability and mode collapse. Flow-based models offer precise likelihood estimation and high-quality generation but require significant computational resources and involve complex model designs. Diffusion models excel in generating high-quality, detailed output, offering flexibility and solid probabilistic foundations, but suffer from long training times and high computational costs. However, many variants and improvements of diffusion models have mitigated these issues, as described in Sec. 2.3.4.
3. Overview and Categorization
@@
Table 1. Summary and main papers of the diffusion models for time series and spatio-temporal data modeling. Red indicates univariate time series, blue-violet is multivariate time series, and yellow shows spatio-temporal data.
@@
WaveGrad (Chen et al., 2020) Univariate DDPM Generation Audio Google ICLR 2021
@@
DiffWave (Kong et al., 2020) Univariate DDPM Generation Audio UCSD ICLR 2021
D-Va (Koa et al., 2023) Univariate DDPM Forecasting Finance NUS CIKM 2023
DiffLoad (Wang et al., 2023b) Univariate DDPM Forecasting Electricity HKU ArXiv 2023
DiffuASR (Liu et al., 2023g) Univariate DDPM Generation Recommendation XJTU CIKM 2023
DiffRec (Wang et al., 2023c) Univariate DDPM Generation Recommendation NUS SIGIR 2023
@@
TimeGrad (Rasul et al., 2021) Multivariate DDPM Forecasting General Zalando ICML 2021
CARD (Han et al., 2022) Multivariate DDPM Classification General UT-Austin NeurIPS 2022
BVAE (Li et al., 2022a) Multivariate DDPM Forecasting General Baidu NeurIPS 2023
@@
SSSD (Alcaraz and Strodthoff, 2022) Multivariate DDPM Imputation General Oldenburg TMLR 2023
DA-TASWDM (Xu et al., 2023c) Multivariate DDPM Imputation Healthcare HKBU CIKM 2023
DiffEEG (Shu et al., 2023) Multivariate DDPM Forecasting Healthcare USTC ArXiv 2023
Tosato et al. (Tosato et al., 2023) Multivariate DDPM Classification Healthcare Tilburg Synapsium 2023
DiffCharge (Li et al., 2024e) Multivariate DDPM Generation Electricity HKUST ArXiv 2023
@@
Yang et al. (Yang et al., 2023a) Multivariate DDPM Imputation&Anomaly Detection AIOps Microsoft ESEC/FSE 2023
@@
AnoDDPM (Sui et al., 2024) Multivariate DDPM Anomaly Detection General Beihang IEEE Sens. J. 2024
DiffShape (Liu et al., 2024b) Multivariate DDPM Classification General SCUT AAAI 2024
@@
STPP (Yuan et al., 2023a) Spatio-temporal DDPM Forecasting General Event THU KDD 2023
@@
ERDiff (Wang et al., 2024b) Spatio-temporal DDPM Alignment General GIT NeurIPS 2023
@@
SpecSTG (Lin et al., 2024) Spatio-temporal DDPM Forecasting Transportation USYD ArXiv 2024
@@
DST-DDPM (Chen et al., 2024b) Spatio-temporal DDPM Forecasting Environment HKUST ENVIRON RES 2024
@@
UTD-PTP (Tang et al., 2024) Spatio-temporal DDPM Forecasting Transportation BIT IEEE Sens. J. 2024
@@
SGMSE (Welker et al., 2022) Univariate Score-based Generation Audio UHH Interspeech 2022
@@
DeScoD-ECG (Li et al., 2023b) Univariate Score-based Denoising Healthcare U of A IEEE J BIOMED HEALTH 2023
ScoreGrad (Yan et al., 2021) Multivariate Score-based Forecasting General BIT ArXiv 2021
Bilo et al. (Biloš et al., 2022) Multivariate Score-based Forecasting General TUM ICML 2023
@@
StoRM (Lemercier et al., 2023) Multivariate Score-based Denoising&Generation Audio UHH IEEE-ACM T AUDIO SPE 2023
@@
Risk-sensitive SDE (Li et al., 2024c) Multivariate Score-based Generation General Cambridge ArXiv 2024
@@
TimeADDM (Hu et al., 2024) Multivariate Score-based Anomaly Detection General HFUT ICASSP 2024
@@
Sasdim (Zhang et al., 2023d) Spatio-temporal Score-based Imputation General CSU ArXiv 2023
@@
Dyffusion (Rühling Cachay et al., 2024) Spatio-temporal Score-based Forecasting General UCSD NeurIPS 2023
Lu et al. (Lu et al., 2022a) Univariate Conditional Generation Audio CMU ICASSP 2022
PulseDiff (Jenkins et al., 2023) Univariate Conditional Imputation Healthcare IC ArXiv 2023
DR-DiffuSE (Tai et al., 2023) Univariate Conditional Generation Audio UESTC AAAI 2023
Dose (Tai et al., 2024) Univariate Conditional Generation Audio UESTC NeurIPS 2023
@@
FUSE (Yang et al., 2024a) Univariate Conditional Generation Audio Oxford Interspeech 2024
@@
DiffsFormer (Gao et al., 2024) Univariate Conditional Imputation&Generation Finance USTC ArXiv 2024
@@
DreamRec (Yang et al., 2024b) Univariate Conditional Generation Recommendation USTC NeurIPS 2024
@@
CSDI (Tashiro et al., 2021) Multivariate Conditional Imputation General Stanford NeurIPS 2021
@@
TimeDiff (Shen and Kwok, 2023) Multivariate Conditional Forecasting General HKUST ICML 2023
@@
DiffAD (Xiao et al., 2023) Multivariate Conditional Anomaly Detection General HENU KDD 2023
@@
D3(Wang et al., 2024c) Multivariate Conditional Anomaly Detection General BUPT NeurIPS 2023
MIDM (Wang et al., 2023d) Multivariate Conditional Imputation General USTC KDD 2023
MEDiC (Sharma et al., 2023) Multivariate Conditional Imputation Healthcare IIT NeurIPS 2023
DiffPLF (Li et al., 2024f) Multivariate Conditional Forecasting Electricity HKUST(GZ) ArXiv 2024
@@
ImDiffusion (Chen et al., 2023c) Multivariate Conditional Anomaly Detection General PKU VLDB 2024
@@
DiffDA (Huang et al., 2024) Multivariate Conditional Forecasting Climate ETH Zurich ArXiv 2024
@@
DiffSTOCK (Daiya et al., 2024) Multivariate Conditional Forecasting Finance Purdue ICASSP 2024
RF-Diffusion (Chi et al., 2024) Multivariate Conditional Generation Network THU MobiCom 2024
@@
DiffSTG (Wen et al., 2023) Spatio-temporal Conditional Forecasting General BJTU SIGSPATIAL 2023
@@
PriSTI (Liu et al., 2023c) Spatio-temporal Conditional Imputation General Beihang ICDE 2023
@@
DiffUFlow (Zheng et al., 2023) Spatio-temporal Conditional Forecasting Transportation CSU CIKM 2023
@@
DiffTraj (Zhu et al., 2024a) Spatio-temporal Conditional Generation Transportation SUSTech NeurIPS 2023
@@
ControlTraj (Zhu et al., 2024b) Spatio-temporal Conditional Generation Transportation HKUST(GZ) arXiv 2024
@@
Diff-RNTraj (Wei et al., 2024) Spatio-temporal Conditional Generation Transportation BJTU ArXiv 2024
Evans et al. (Evans et al., 2024) Univariate LDM Generation Audio Stability AI ArXiv 2024
LDT (Feng et al., 2024) Multivariate LDM Forecasting General NTU AAAI 2024
Aristimunha et al. (Aristimunha et al., 2023) Multivariate LDM Generation Healthcare UPSaclay NeurIPS 2023
@@
TSDM (Pei et al., 2023) Multivariate DDIM Imputation&Anomaly Detection Electricity HUST ArXiv 2023
@@
LADM (Lv et al., 2024) Spatio-temporal LDM Forecasting Transportation XJU IEEE T INSTRUM MEAS 2024
Diffusion Models for